OpenStudy (anonymous):

The tram at Valley View Ski Resort can transport 5000 people in 6 hours. Using the chairlift and tram, 5000 people can be brought to the top in 2 hours. If the tram breaks, how long will it take the chairlift to bring 5000 to the mountaintop? @Calculus1

OpenStudy (anonymous):

what the heck kind of problem is this?

OpenStudy (anonymous):

I believe its math, but my friends need help on this and I'm not sure how to solve it.

OpenStudy (anonymous):

i mean we can do it, but it is not your usual calc kind of problem. just need the rate for each

OpenStudy (anonymous):

i think its a work rate problem...

OpenStudy (anonymous):

yeah, my friend is in a non-cal math then me, lol. Right now in my calculus class were doing sigma notation

OpenStudy (anonymous):

i think the answer is \[\frac{11}{6}\] but i am not confident so let me check

OpenStudy (anonymous):

could you show the work on how to work

OpenStudy (anonymous):

the problem

OpenStudy (anonymous):

ok let me see if i can write something that make sense. @simon if you have a better explanation please jump in

OpenStudy (anonymous):

first of all forget the 5000, because that is a red herring. it is 5000 all the way across, so just think of it as 1, as in one job

OpenStudy (anonymous):

@satellite73 I got a different answer: 3 hours

OpenStudy (anonymous):

you know the rate of the tram is \[\frac{1}{6}\] because it carries one sixth of the people per hour. so lets put the rate of the chair lift as R

OpenStudy (anonymous):

yeah i think i was wrong

OpenStudy (anonymous):

so their combined rate is \[\frac{1}{6}+R\] and you also know that it takes 2 hours to complete one job, so you know \[(\frac{1}{6}+R)\times 2=1\] and now we can solve for R

OpenStudy (anonymous):

rate of what? I thought of it as the rate\[\frac{dP}{dt} = \frac{5000}{6}\]

OpenStudy (anonymous):

I think it's not a good idea to forget about the number 5000

OpenStudy (anonymous):

and so you get just what victorana said, \[R=\frac{1}{3}\] meaning it does one third of the job in an hour, or one job in 3 hours

OpenStudy (anonymous):

you can work with the 5000 all the way across if you want, but it is 5000 in every spot. so just make the 5000 a 1 to compute. still get the same thing

OpenStudy (anonymous):

just like all those problems that say "bill can paint the house if 4 hours, joe can paint it in 6 hours, how long if they work together?" just make the job a 1 and forget about how many square feet they are painting

OpenStudy (anonymous):

Well I'm not pretty sure that your idea it's ok, but it solves the problem. I have to think!

OpenStudy (anonymous):

I got it, that was very clear! Thanks!

OpenStudy (anonymous):

i am just replacing 5000 by "all the people"

OpenStudy (anonymous):

yw, sorry it took a while

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw, sorry if the explanation was not so clear, i got off on a false start. forgot that i should set \[(\frac{1}{6}+R)\times 2 = 1\]